OpenStudy (anonymous):

What is the solution of 2(x + 4) - x = 15? or 2x + 8 -x = 15

myininaya (myininaya):

@Tea-Is-God Try to distribute first 2(x+4)

OpenStudy (anonymous):

2x + 4 - x = 15 can you figure it out from there

myininaya (myininaya):

Well @moser90 Your distribution is a little off

myininaya (myininaya):

Because 2(x+4)=2(x)+2(4)=2x+8

OpenStudy (anonymous):

oops plus 8

myininaya (myininaya):

2x+8-x=15 Now you have like terms on the left hand side

myininaya (myininaya):

Go ahead and perform the operation that combines them

OpenStudy (anonymous):

3x=7?

OpenStudy (anonymous):

x=7?

myininaya (myininaya):

2x-x+8=15 2x-x=15-8 yes!

OpenStudy (anonymous):

so its 7=x